OpenStudy (anonymous):

Which equation would best help solve the following problem? Hideki throws a baseball ball upward with an initial velocity of 12 m/s. When he throws the ball, it is 2 m above the ground. How long will it take the ball to hit the ground? a. –16t2+ 12t + 2 = 0 b. –4.9t2 + 12t + 2 = 0 c. –4.9t2 + 12t = 2 d. 4.9t2 + 12t + 2 = 0

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

@Michele_Laino

jimthompson5910 (jim_thompson5910):

The general equation is \[\Large h = -\frac{g}{2}t^2 + v*t + s\] h = height at time t g = acceleration of gravity (g = 9.8 m/s^2 approximately) v = initial velocity s = initial height

OpenStudy (anonymous):

I need an example of this.... I don't understand

jimthompson5910 (jim_thompson5910):

ok let's say initial velocity is v = 10 m/s initial height is s = 12 m that would mean the height equation is \[\Large h = -\frac{g}{2}t^2 + v*t + s\] \[\Large h = -\frac{9.8}{2}t^2 + 10t + 12\] \[\Large h = -4.9t^2 + 10t + 12\]

jimthompson5910 (jim_thompson5910):

solving h(t) = 0 will give you the time values when the object is on the ground